a. It is said that in B-trees with realistically large minimum degrees, almost all keys are stored in the leaf nodes. Justify this statement mathematically. Numerically demonstrate your mathematical justification with examples of t = 100 and 200.

b. Modify B-Tree-Search(x, k) so that it returns all occurrences of k in the tree pointed to by x. The modified function is to return the set of all (p, i) where p is a pointer to a node containing keyi = k.
Answer #1

a)

If minimum degree is large and each node have maximum number of nodes, then the B-tree will have minimum height and almost all the keys are stored in the leaf nodes.

Assume height of the B-tree is h and each node in the B-tree is full. Then the total number of keys in the tree , n= (2t)h+1-1.

The number of keys at the depth h or leaves in the B-tree= [(2t)h+1-1]-[ (2t)h-1]

The ratio between total number of keys and the number of leaves( or percentage of leaves in the total number of keys)

= [(2t)h+1-1]-[ (2t)h-1] / (2t)h+1-1 ≈ 1-1/2t= (2t-1)/2t > 99%

Therefore, almost keys are stored in the leaves.

For suppose , t=100 then (2t-1)/2t =200-1/200=199/200= 0.995 or 99.5%.

Thus, if t=100, 99.5% of keys will be stored in the leaves.

For suppose , t=200 then (2t-1)/2t =400-1/400=399/400= 0.9975 or 99.75%.

Thus, if t=200, 99.75% of keys will be stored in the leaves.

b)

The above modified B-TREE-SEARCH takes a root node and a key k to be searched as input and returns the all occurrences of k in the tree pointed to by x.